# thanks: ORCID ID: 0009-0003-9679-035Xthanks: ORCID ID: $0$-0001-7989-9271

### v1

@@ Appendix A Dynamical equations in hydrodynamic form
$\mu_{\alpha}$ $=\frac{\delta F}{\delta\tilde{\Gamma}_{\alpha}}=\partial_{\Gamma_{\alpha}}g\hskip 14.22636pt\text{with}\hskip 14.22636pt\alpha=1,2.$
@@
$p$ $=\partial_{h}f-\nabla\cdot\left[\frac{1}{\xi}\left(\gamma_{0}-k_{b}T\Gamma_{1}-k_{b}T\Gamma_{2}\right)\nabla h\right]$ (17)
@@
Here, $\gamma(\Gamma_{1},\Gamma_{2})=\gamma_{0}-k_{b}T(\Gamma_{1}+\Gamma_{2})$ is the resulting local surface tension. By substituting the expression for the chemical potentials in Eqs. (17) into Eqs. (7) and transitioning to the long-wave limit $|\nabla h|\ll 1$ (yielding $\tilde{\Gamma}_{\alpha}\approx\Gamma_{\alpha}$), one obtains the hydrodynamic form
@@ Appendix B Nondimensionalization
We rescale Eqs. (18) by introducing the nondimensionalization
@@ Appendix C Hopf bifurcations of the local reactor
$0=r\left[2\delta\Gamma_{2,ss}\Gamma_{1,ss}^{2}-(3\delta^{2}+\delta)\Gamma_{1,ss}^{2}\right]-\beta_{1}\delta.$ (27)
@@
$\frac{\partial\left(\Gamma_{1,ss}+\Gamma_{2,ss}\right)}{\partial\Gamma_{1,ss}}=0.$ (29)
By using Eqs. (12) one also obtains the relation $\Gamma_{2,ss}=\delta^{2}\Gamma_{1,ss}+\frac{1}{r\delta\Gamma_{1,ss}^{2}}\left(\delta\beta_{1}\Gamma-\beta_{1}e^{\mu_{1}}\right)$. Substituting this into (29) again yields (28).
@@ Appendix D Approximations for flux-reaction balance
00 0 $=-\partial_{z}j_{1}+\mathcal{R}+\mathcal{B}_{1}-v\partial_{z}\Gamma_{1},$ (30)
00 0 $=-\partial_{z}j_{2}-\mathcal{R}+\mathcal{B}_{2}-v\partial_{z}\Gamma_{2}.$
@@
First, we have assumed that the extrema of $\Gamma_{1},\Gamma_{2}$ and $\gamma$ coincide in space. This allows us to neglect the terms $-v\partial_{z}\Gamma_{1},-v\partial_{z}\Gamma_{2}$ at the surface tension peak ($z=z_{m}$) in region (ii) [see Fig. 14(a)], yielding the balance equations
00 0 $=-\partial_{z}j_{1}+\mathcal{R}+\mathcal{B}_{1},$ (31)
00 0 $=-\partial_{z}j_{2}-\mathcal{R}+\mathcal{B}_{2},$
@@
at $z=z_{m}$ in the comoving frame. Second, we have argued that the transport contribution $-\partial_{z}j_{1}$ is negligible at the peak since diffusion and advective contributions effectively cancel. One then obtains
00 0 $=\mathcal{R}+\mathcal{B}_{1},$ (32)
00 0 $=-\partial_{z}j_{2}-\mathcal{R}+\mathcal{B}_{2},$
@@
As described in the main text, Eqs. (32) correspond to the steady state equations of the local reactor with the shifted parameter $\beta_{2}e^{\mu_{2}}-\partial_{z}j_{2}(z_{m})$, i.e., the surface tension at $z_{m}$ can be obtained by evaluating $\gamma_{ss}(\beta_{2}e^{\mu_{2}}-\partial_{z}j_{2}(z_{m}))$. Analogously, if $-\partial_{z}j_{1}$ is not neglected, one may obtain the peak surface tension by shifting both driving parameters, i.e., by evaluating $\gamma_{ss}(\beta_{2}e^{\mu_{2}}-\partial_{z}j_{2}(z_{m}),\beta_{1}e^{\mu_{1}}-\partial_{z}j_{1}(z_{m}))$. For the local reactor, this corresponds to considering the function $\gamma_{ss}(\beta_{2}e^{\mu_{2}},\beta_{1}e^{\mu_{1}})$ instead, where $\beta_{1}e^{\mu_{1}}$ is a secondary parameter [cf. Fig. 3(f)]. We generally find that Eqs. (32) are appropriate to describe $\gamma_{2}$, i.e., it indeed holds that $\gamma_{2}=\gamma_{ss}(\beta_{2}e^{\mu_{2}}-\partial_{z}j_{2}(z_{m}))$ [Figs. 4(b) and 14(b)]. However, a caveat is in order. By comparing the steady state curves $\gamma_{ss}$ either with or without shifting $\beta_{1}e^{\mu_{1}}$ (the steady state surface tension corresponding to either Eqs. (31) or (32), respectively) we find that these curves differ significantly, i.e., the contribution $-\partial_{z}j_{1}(z_{m})$ is not negligible. The reason is that the terms $-v\partial_{z}\Gamma_{1},-v\partial_{z}\Gamma_{2}$ and $-\partial_{z}j_{1}$ do not vanish at the same point in space (if they did, our approximations would be exact) and the corresponding contributions in Eqs. (30) are large enough to significantly change $\gamma_{ss}$ when absorbed into the parameters of the local reactor. The circumstance that $\gamma_{2}$ is nevertheless well captured by Eqs. (32) is in fact a consequence of a favorable cancellation of the errors induced by transitioning from Eqs. (30) to (31) and finally to (32). However, we find that in practice [for diffusion coefficients ${D=D_{1}=D_{2}=\mathcal{O}(10^{-1})}$] one can always find a point $z_{m}^{*}$ in the immediate vicinity of $z_{m}$ with comparable surface tension $\gamma_{2}^{*}$ where all three terms $-v\partial_{z}\Gamma_{1}(z_{m}^{*}),-v\partial_{z}\Gamma_{2}(z_{m}^{*})$ and $-\partial_{z}j_{1}(z_{m}^{*})$ are indeed negligible (Fig. 14). Therefore, Eqs. (32) generally hold for some point near the surface tension peak with a surface tension that is comparable to the maximum value. This point is usually not exactly the surface tension maximum. Because the differences $\Delta\gamma=\gamma_{2}-\gamma_{1}$ and $\Delta\gamma^{*}=\gamma_{2}^{*}-\gamma_{1}$ are nearly identical, we may then draw the same conclusions regarding the interplay between the Marangoni effect and the local reactor as in Sec. 3.
@@
Figure 14: (a) Magnification of the advancing contact line region of a self-propelled droplet. The top panel shows the film-height profile and the velocity field of the liquid, the bottom panel shows the surface tension profile. The two points $z_{m}$ and $z_{m}^{*}$ are the surface tension maximum ($\gamma=\gamma_{2}$) and a point slightly offset from it ($\gamma=\gamma_{2}^{*}$). (b) Steady state curve $\gamma_{ss}$ of the local reactor as a function of $\beta_{2}e^{\mu_{2}}-\partial_{z}j_{2}$. We consider three values of the secondary parameter (see text), corresponding to $\beta_{1}e^{\mu_{1}}$ (solid orange line), $\beta_{1}e^{\mu_{1}}-\partial_{z}j_{1}(z_{m})$ (dashed black line) and $\beta_{1}e^{\mu_{1}}-\partial_{z}j_{1}(z_{m}^{*})$ (dashed gray line). When considering $\gamma_{2}$, there are significant deviations between the cases $\beta_{1}e^{\mu_{1}}$ and $\beta_{1}e^{\mu_{1}}-\partial_{z}j_{1}(z_{m})$. For $\gamma_{2}^{*}$, the cases $\beta_{1}e^{\mu_{1}}$ and $\beta_{1}e^{\mu_{1}}-\partial_{z}j_{1}(z_{m}^{*})$ nearly overlap and $\gamma_{2}^{*}$ approximately lies on both curves.
@@ Appendix E Numerical Methods
Direct numerical simulations are based on the finite-element method and are implemented using the open source library oomph-lib [65]. Most simulations are performed on a periodic domain that is discretized using a static spatial mesh (1D: 641 nodes, 2D: $251\times 251$ nodes). An exception are the data shown in Figs. 8(a)-8(c), where we use an adaptive spatial mesh. For temporal discretization, we use a backward differentiation scheme of order 2 (BDF(2)) with adaptive time stepping.
For numerical continuation, Eqs. (23) are spatially discretized using the finite element method on a one-dimensional periodic domain (400 nodes). We employ the continuation package pde2path [69] which uses pseudo-arc length continuation with a predictor-corrector method. Because Eqs. (23) exhibit continuous symmetries (liquid volume conservation, translational symmetry on a periodic domain), they are augmented by algebraic constraints for the liquid volume and the spatial phase. Then, two additional parameters must be freed which are also determined during continuation. In particular, the parameter corresponding to the spatial phase constraint is the velocity $v$ of the comoving frame, i.e., the droplet speed [Fig. 5(d)]. For details, see Ref. [69].
@@ Acknowledgements
Part of the calculations for this publication were performed on the HPC cluster PALMA II of the University of Münster, subsidized by the Deutsche Forschungsgemeinschaft (DFG) (INST 211/667-1). We acknowledge financial support by the DPG via the grant no. TH 781/12-2 within SPP 2171. F.V. further acknowledges valuable discussions with Yutaka Sumino and wishes to thank the entire Sumino Lab for their hospitality during his stay in Japan. UT would like to thank the Kavli Institute for Theoretical Physics (KITP), Santa Barbara (grant NSF PHY-2309135) for support and hospitality during the programme Active Solids where part of the work was undertaken.
@@
The data and python codes for creating the figures for this study will be made publicly available at zenodo.

### v3

@@ Appendix A Dynamical equations in hydrodynamic form
$\mu_{\alpha}$ $=\frac{\delta F}{\delta\hat{\Gamma}_{\alpha}}=\partial_{\Gamma_{\alpha}}g\hskip 14.22636pt\text{with}\hskip 14.22636pt\alpha=1,2.$
@@
$p$ $=\partial_{h}f-\nabla\cdot\left[\frac{1}{\xi}\gamma(\Gamma_{1},\Gamma_{2})\nabla h\right]$ (17)
@@
Here, $\gamma(\Gamma_{1},\Gamma_{2})=\gamma_{0}-k_{b}T(\Gamma_{1}+\Gamma_{2})$ is the resulting local surface tension. By substituting the expression for the chemical potentials in Eqs. (17) into Eqs. (7) and transitioning to the long-wave limit $|\nabla h|\ll 1$ (yielding $\hat{\Gamma}_{\alpha}\approx\Gamma_{\alpha}$), one obtains the hydrodynamic form
@@ Appendix B Nondimensionalization
We rescale Eqs. (18) by introducing the scalings
@@
[…]
Table 1: Example set of approximate numerical values for the scales employed in the nondimensionalization (20) and the dimensional parameters of the model (7).
Symbol Quantity Numerical Value
[…]
$a_{1},a_{2}$ surfactant length scales $3\text{\,}\frac{\mathrm{nm}}{}$
[…]
$\eta$ dynamic viscosity $0.2\text{\,}\frac{\mathrm{Pa}\text{\,}\mathrm{s}}{}$
$\gamma_{0}$ surface tension (bare droplet) $70\text{\times}{10}^{-3}\text{\,}\frac{\mathrm{J}}{{\mathrm{m}}^{2}}$
$\kappa$ thermal energy density of surfactant $4\text{\times}{10}^{-4}\text{\,}\frac{\mathrm{J}}{{\mathrm{m}}^{2}}$
[…]
$\tau$ time scale $0.03\text{\,}\frac{\mathrm{s}}{}$
AA Hamaker constant $1\text{\times}{10}^{-15}\text{\,}\frac{\mathrm{J}}{}$
$\theta_{eq}$ equilibrium contact angle (bare droplet) $10\text{\,}\frac{\mathrm{\SIUnitSymbolDegree}}{}$
$\frac{L^{2}}{\tau}\tilde{D}_{1},\frac{L^{2}}{\tau}\tilde{D}_{2}$ surfactant diffusivity (on interface) $4\text{\times}{10}^{-10}\text{\,}\frac{{\mathrm{m}}^{2}}{\mathrm{s}}$
$\frac{1}{\tau}\tilde{r},\frac{1}{\tau}\tilde{\beta}_{1},\frac{1}{\tau}\tilde{\beta}_{2}$ reaction rate constants $10\text{\,}\frac{}{\mathrm{s}}$, $0.3\text{\,}\frac{}{\mathrm{s}}$$70\text{\,}\frac{}{\mathrm{s}}$
$\mu_{1},\mu_{2}$ chemical potentials (chemostats) -1 $k_{b}T$, 4 $k_{b}T$
@@ Appendix C Hopf bifurcations of the local reactor
$0=r\left[2\delta\Gamma_{2,ss}\Gamma_{1,ss}-(3\delta^{3}+\delta)\Gamma_{1,ss}^{2}\right]-\delta\beta_{1}.$ (27)
@@
$\frac{\partial\left(\Gamma_{1,ss}+\Gamma_{2,ss}\right)}{\partial\Gamma_{1,ss}}=0,$ (29)
where $\Gamma_{2,ss}$ is understood as a function of $\Gamma_{1,ss}$ and of all other parameters (except $\beta_{2}$ and $\mu_{2}$) appearing in Eqs. (12). By using Eqs. (12) one also obtains the relation $\Gamma_{2,ss}=\delta^{2}\Gamma_{1,ss}+\frac{1}{r\delta\Gamma_{1,ss}^{2}}\left(\delta\beta_{1}\Gamma_{1,ss}-\beta_{1}e^{\mu_{1}}\right)$. Substituting this into (29) again yields (28).
@@ Appendix D Approximations for flux-reaction balance
$0$ $=-\partial_{z}j_{1}+\mathcal{R}+\mathcal{B}_{1}-v\partial_{z}\Gamma_{1},$ (30)
$0$ $=-\partial_{z}j_{2}-\mathcal{R}+\mathcal{B}_{2}-v\partial_{z}\Gamma_{2}.$
@@
First, we have assumed that the extrema of $\Gamma_{1},\Gamma_{2}$ and $\gamma$ coincide in space. This allows us to neglect the terms $v\partial_{z}\Gamma_{1},-v\partial_{z}\Gamma_{2}$ at the surface tension peak ($z=z_{m}$) in region (ii) [see Fig. 14(a)], yielding the balance equations
$0$ $=-\partial_{z}j_{1}+\mathcal{R}+\mathcal{B}_{1},$ (31)
$0$ $=-\partial_{z}j_{2}-\mathcal{R}+\mathcal{B}_{2},$
@@
at $z=z_{m}$ in the comoving frame. Second, we have argued that the transport contribution $\partial_{z}j_{1}$ is negligible at the peak since diffusive and advective contributions effectively cancel. One then obtains
$0$ $=\mathcal{R}+\mathcal{B}_{1},$ (32)
$0$ $=-\partial_{z}j_{2}-\mathcal{R}+\mathcal{B}_{2},$
As described in the main text, Eqs. (32) correspond to the steady state equations of the local reactor with the shifted parameter $\beta_{2}e^{\mu_{2}}-\partial_{z}j_{2}(z_{m})$, i.e., the surface tension at $z_{m}$ can be obtained by evaluating $\gamma_{ss}(\beta_{2}e^{\mu_{2}}-\partial_{z}j_{2}(z_{m}))$. Analogously, if $\partial_{z}j_{1}$ is not neglected, one may obtain the peak surface tension by shifting both driving parameters, i.e., by evaluating $\gamma_{ss}(\beta_{2}e^{\mu_{2}}-\partial_{z}j_{2}(z_{m}),\beta_{1}e^{\mu_{1}}-\partial_{z}j_{1}(z_{m}))$. For the local reactor, this corresponds to considering the function $\gamma_{ss}(\beta_{2}e^{\mu_{2}},\beta_{1}e^{\mu_{1}})$ instead, where $\beta_{1}e^{\mu_{1}}$ is a secondary parameter [cf. Fig. 3(f)]. Although we justify dropping the contribution $\partial_{z}j_{1}(z_{m})$ in the following, we point out that doing so is not strictly necessary. Rather, it simplifies the discussion in Sec. 3. From Fig. 3(f) it can be seen that changes in $\beta_{1}e^{\mu_{1}}$ [here $\beta_{1}e^{\mu_{1}}-\partial_{z}j_{1}$] only influence the shape of $\gamma_{ss}$ as a function of $\beta_{2}e^{\mu_{2}}$ [here $\beta_{2}e^{\mu_{2}}-\partial_{z}j_{2}$] near its local minimum, while the curve remains mostly unaffected for larger values of $\beta_{2}e^{\mu_{2}}$ near its local maximum. It is this section of $\gamma_{ss}$ near its local maximum that is relevant for points in space near the localized surface tension peak at $z=z_{m}$. Therefore, using Eq. (31) instead of (32) only marginally affects $\gamma_{2}$.
@@
Nevertheless, we generally find that Eqs. (32) are appropriate to describe $\gamma_{2}$, i.e., it indeed holds that $\gamma_{2}=\gamma_{ss}(\beta_{2}e^{\mu_{2}}-\partial_{z}j_{2}(z_{m}))$ [Figs. 4(b) and 14(b)]. However, a caveat is in order. By comparing the curves $\gamma_{ss}$ as a function of $\beta_{2}e^{\mu_{2}}-\partial_{z}j_{2}$ either with or without shifting $\beta_{1}e^{\mu_{1}}$ (the steady state surface tension corresponding to either Eqs. (31) or (32), respectively) we find that these curves differ significantly, i.e., the contribution $\partial_{z}j_{1}(z_{m})$ is not negligible [Fig. 14(b)]. The reason is that the terms $v\partial_{z}\Gamma_{1},-v\partial_{z}\Gamma_{2}$ and $\partial_{z}j_{1}$ do not vanish at the same point in space (if they did, our approximations would be exact) and the corresponding contributions in Eqs. (30) are large enough to significantly change $\gamma_{ss}$ when absorbed into the parameters of the local reactor. The circumstance that $\gamma_{2}$ is nevertheless well captured by Eqs. (32) is in fact a consequence of a favorable cancellation of the errors induced by transitioning from Eqs. (30) to (31) and finally to (32). However, we find that in practice [for diffusion coefficients $D=D_{1}=D_{2}=\mathcal{O}(10^{-1})$}] one can always find a point $z_{m}^{*}$ in the immediate vicinity of $z_{m}$ with comparable surface tension $\gamma_{2}^{*}$ where all three terms $v\partial_{z}\Gamma_{1}(z_{m}^{*}),-v\partial_{z}\Gamma_{2}(z_{m}^{*})$ and $\partial_{z}j_{1}(z_{m}^{*})$ are indeed negligible (Fig. 14). Therefore, the Eqs. (32) generally hold for some point near the surface tension peak with a surface tension that is comparable to the maximum value. This point is usually not exactly the surface tension maximum. Because the differences $\Delta\gamma=\gamma_{2}-\gamma_{1}$ and $\Delta\gamma^{*}=\gamma_{2}^{*}-\gamma_{1}$ are nearly identical, we may then draw the same conclusions regarding the interplay between the Marangoni effect and the local reactor as in Sec. 3.
@@
Figure 14: (a) Magnification of the advancing contact line region of a self-propelled droplet in the comoving frame z=xvtz=x-vt. The top panel shows the film-height profile and the velocity field of the liquid (in the laboratory frame), the bottom panel shows the surface tension profile. The two points $z_{m}$ and $z_{m}^{*}$ are the surface tension maximum ($\gamma=\gamma_{2}$) and a point slightly offset from it ($\gamma=\gamma_{2}^{*}$). (b) Steady state curve $\gamma_{ss}$ of the local reactor as a function of $\beta_{2}e^{\mu_{2}}-\partial_{z}j_{2}$. We consider three values of the secondary parameter (see text), corresponding to $\beta_{1}e^{\mu_{1}}$ (solid orange line), $\beta_{1}e^{\mu_{1}}-\partial_{z}j_{1}(z_{m})$ (dashed black line) and $\beta_{1}e^{\mu_{1}}-\partial_{z}j_{1}(z_{m}^{*})$ (dashed green line). When considering $\gamma_{2}$, there are significant deviations between the cases $\beta_{1}e^{\mu_{1}}$ and $\beta_{1}e^{\mu_{1}}-\partial_{z}j_{1}(z_{m})$. For $\gamma_{2}^{*}$, the cases $\beta_{1}e^{\mu_{1}}$ and $\beta_{1}e^{\mu_{1}}-\partial_{z}j_{1}(z_{m}^{*})$ nearly overlap and $\gamma_{2}^{*}$ approximately lies on both curves.
@@ Appendix E Numerical Methods
Direct numerical simulations are based on the finite-element method and are implemented using the open source library oomph-lib [75]. Most simulations are performed on a periodic domain that is discretized using a static spatial mesh (1D: 641 nodes, 2D: $251\times 251$ nodes). An exception are the data shown in Figs. 4(a), 8(a),8(c), 10(a) and 14(a), where we use an adaptive spatial mesh. For temporal discretization, we use a backward differentiation scheme of order 2 (BDF(2)) with adaptive time stepping.
For numerical continuation [77, 78, 79, 80], Eqs. (23) are spatially discretized using the finite element method on a one-dimensional periodic domain (400 nodes). We employ the continuation package pde2path [80] which uses pseudo-arc length continuation with a predictor-corrector method. Because Eqs. (23) exhibit continuous symmetries (liquid volume conservation, translational symmetry on a periodic domain), they are augmented by algebraic constraints for the liquid volume and the spatial phase. Then, two additional parameters must be freed which are also determined during continuation. In particular, the parameter corresponding to the spatial phase constraint is the velocity $v$ of the comoving frame, i.e., the droplet speed [Fig. 5(d)]. For details, see Refs. [79, 80].
@@ Acknowledgements
Part of the calculations for this publication were performed on the HPC cluster PALMA II of the University of Münster, subsidized by the Deutsche Forschungsgemeinschaft (DFG) (INST 211/667-1). We acknowledge financial support by the DPG via the grant no. TH 781/12-2 within SPP 2171. FV further acknowledges valuable discussions with Yutaka Sumino and wishes to thank the entire Sumino Lab for their hospitality during his stay in Japan. UT would like to thank the Kavli Institute for Theoretical Physics (KITP), Santa Barbara (grant NSF PHY-2309135) for support and hospitality during the programme Active Solids where part of the work was undertaken.
@@
The data and python codes used in this study as well as the supplemental videos referenced in the main text are publicly available on Zenodo [124].